A sample of 6 rivets is randomly selected and the following diameters are measured in millimeters. Give a point estimate for the population variance. Round your answer to three decimal places.

8.01, 7.99, 8.05, 8.00, 8.04, 8.05
Answer #1

Solution :

n = 6

= = (8.01 + 7.99 + 8.05 + 8.00 + 8.04 + 8.05) / 6 = 8.023

A point estimate for the population variance is sample variance .

s2 =  (x - )2 / n - 1  

= 0.0035333 / 5

= 0.001